# TabPFN: One Model to Rule Them All?

### v1

@@ Abstract
Hollmann et al. (Nature 637 (2025) 319–326) recently introduced TabPFN, a transformer-based deep learning model for regression and classification on tabular data, which they claim “outperforms all previous methods on datasets with up to 10,000 samples by a wide margin, using substantially less training time.” Furthermore, they have called TabPFN a “foundation model” for tabular data, as it can support “data generation, density estimation, learning reusable embeddings and fine-tuning”. If these statements are well-supported, TabPFN may have the potential to supersede existing modeling approaches on a wide range of statistical tasks, mirroring a similar revolution in other areas of artificial intelligence that began with the advent of large language models. In this paper, we provide a tailored explanation of how TabPFN works for a statistics audience, by emphasizing its interpretation as approximate Bayesian inference. We also provide more evidence of TabPFN’s “foundation model” capabilities: We show that an out-of-the-box application of TabPFN vastly outperforms specialized state-of-the-art methods for semi-supervised parameter estimation, prediction under covariate shift, and heterogeneous treatment effect estimation. We further show that TabPFN can outperform LASSO at sparse regression and can break a robustness-efficiency trade-off in classification. All experiments can be reproduced using the code provided at https://github.com/qinglong-tian/tabpfn_study.
Foundation models,
in-context learning,
meta-learning,
tabular data,
@@ 1 Introduction
Deep learning has achieved remarkable success in modeling unstructured data types such as text and images. However, for tabular data—arguably the most prevalent and consequential data type across science—ensemble tree methods have remained the gold standard predictive modeling strategy, consistently outperforming even deep learning on benchmark datasets [9, 13, 29, 17]. This long-held wisdom has now been challenged by Hollmann et al. [20], who introduced a deep learning method for tabular data, called TabPFN, which they claimed “outperforms all previous methods on datasets with up to 10,000 samples by a wide margin, using substantially less training time.” As evidence, they showed that TabPFN significantly outperforms random forests, XGBoost, LightGBM, and CatBoost on $50$+ benchmark classification and regression datasets.
@@
Beyond its impressive predictive accuracy, TabPFN is claimed to be robust to outliers and missing data, and automatically estimates predictive distributions, thereby providing uncertainty quantification. Furthermore, it can support “data generation, density estimation, learning reusable embeddings and fine-tuning”. Because of this wide range of capabilities, Hollmann et al. [20] call TabPFN a “tabular foundation model” [39], appropriating a term that has been used to describe the transformative role played by large language models (LLMs) [6] in natural language processing. TabPFN shares other similarities with LLMs: It makes use of the same type of deep learning architecture (transformers) [40] and the same learning paradigm (in-context learning) [7], which differs fundamentally from traditional statistical and machine learning approaches.
Given TabPFN’s promising capabilities, the goal of this paper is twofold.
  • First, we aim to provide an explanation of how TabPFN works that is tailored to a statistics audience (Section 2).
  • Second, and more importantly, we aim to explore the significance of TabPFN to the field of statistics.
@@
Towards the second goal, we demonstrate the utility of TabPFN in handling three currently researched estimation problems, namely semi-supervised parameter estimation (Section 3.1), prediction under covariate shift (Section 3.3), and heterogeneous treatment effects estimation (Section 3.2). For each problem, we compare a simple out-of-the-box application of TabPFN to a “state-of-the-art” method proposed by a selected paper published recently in a top statistics journal. Using the same simulation set-up described by the paper, we show that TabPFN achieves significantly better estimation accuracy than the comparator method.
Apart from examining the performance of TabPFN on specialized estimation tasks, we also examine how it adapts to different types of structure in regression and classification problems. We observe that it outperforms parametric methods (LASSO and linear discriminant analysis) when the model is correctly specified, as well as local averaging and smoothing methods when the model is misspecified (such as via label noise in classification). We discuss these and other discoveries regarding the inductive biases of TabPFN in Section 4.
It is not the point of our paper, however, to suggest that TabPFN is a perfect model. Rather, our argument is that the approach it takes—“in-context learning” using a “foundation model”—has the potential to supersede existing modeling approaches on a wide range of statistical tasks, mirroring a similar revolution in other areas of artificial intelligence that began with the advent of large language models (LLMs). We define these key terms and discuss them further in Section 5. We also discuss opportunities for statisticians to influence the field’s development.
@@ 2 An overview of TabPFN
TabPFN makes use of in-context learning (ICL), which is a new learning paradigm that differs from the standard statistical learning approach to regression and classification and which has gathered significant interest in the machine learning community. The meaning of ICL has changed over time. In TabPFN, ICL takes the form of approximate Bayesian inference; we elaborate on this perspective below and defer a discussion on the historical development of ICL to Section 5.
The first key idea of TabPFN is to approximate posterior predictive distributions (PPD) using a transformer model, which is the same deep learning architecture that powers LLMs. More precisely, let $\Pi$ denote the space of joint distributions on covariate-label pairs $(X,Y)$ and let $n$ be a fixed sample size. Given a prior $p$ on $\Pi$ and an observed sequence of training data $\mathcal{D}=(X_{1},Y_{1},\ldots,X_{n},Y_{n})$ assumed to be drawn independently identically distributed (IID) from some fixed distribution $\pi^{*}\in\Pi$, let $p(\pi|\mathcal{D})$ denote the posterior on $\Pi$ given $\mathcal{D}$. With slight abuse of notation, the posterior predictive distribution (PPD) for $Y_{n+1}$ given $X_{n+1}=x$ is defined as
[…]
@@
Approximating the PPD of a Bayesian regression or classification model is far from a new idea. Indeed, the popular Bayesian Additive Regression Trees (BART) algorithm does precisely this via Markov chain Monte Carlo (MCMC) [11]. The novelty of the PFN approach is that while previous approaches (MCMC, variational inference, likelihood-free inference, etc.) attempt to approximate the map $x\mapsto p(y|x,\mathcal{D})$ for a fixed training dataset $\mathcal{D}$, PFN tries to approximate the map jointly over $x$ and $\mathcal{D}$. It has the computational benefit of amortizing computation, which is the basis of the “substantially less training time” claim by Hollmann et al. [20] regarding TabPFN. [^1] Similar ideas have been explored in the variational inference community under the name of Neural Processes (NP) [16]. Both the NP and PFN approaches use a “meta-learning” perspective to approximate the PPD, but they differ in their implementation choices. NPs employ an encoder-decoder architecture with a parametric distribution over the latent variable $z$, naturally lending itself to variational inference. PFNs, on the other hand, adopt a transformer architecture that avoids explicit distributional assumptions in its internal representations. As a result, their training objectives differ slightly: NPs explicitly optimize a variational lower bound, while PFNs frame learning as a direct optimization problem. Such an approximation may also produce statistical benefits via implicit regularization that are yet to be fully understood.
@@
The second key idea of TabPFN is its choice of prior on the space of joint distribution $\Pi$. The large number of parameters of a transformer model ($\approx$ 7M in the case of TabPFN) means that it could potentially approximate PPDs defined in terms of complicated priors. TabPFN chooses to define its prior in terms of structural causal models (SCMs) [30]. An SCM defines a directed acyclic graph $\mathcal{G}$ whose nodes represent variables $Z_{1},\ldots,Z_{p}$ and whose edges represent causal relationships. Specifically, each variable $Z_{k}$ is defined via the equation $Z_{k}=f_{k}(Z_{\operatorname{PA}_{\mathcal{G}}(k)},\varepsilon_{k})$, where $f_{k}$ is a deterministic function, $\operatorname{PA}_{\mathcal{G}}(k)$ denotes the parents of $Z_{k}$, and $\varepsilon_{1},\ldots,\varepsilon_{p}$ are independent “exogenous” random variables. SCMs have long been used to model causal inference from observational data, in which case $f_{1},\ldots,f_{p}$ are usually defined to be linear functions. Instead, TabPFN allows them to take the form of neural networks or decision trees, thereby modeling nonlinear and even discontinuous structure.
@@
These two key ideas cohere to provide an efficient training strategy for TabPFN. To fit the model’s parameters $\theta$, a collection $\mathfrak{D}$ of approximately 130M synthetic datasets were generated. Each dataset $\mathcal{D}\in\mathfrak{D}$ contains IID examples from a joint distribution $\pi$, which is itself drawn uniformly from the SCM prior. Note that while $\pi$ may have a complex density, the SCM structure gives a simple way to sample from it. Next, $\mathcal{D}$ is partitioned into a “pseudo”-training set $\mathcal{D}_{\text{train}}$ and “pseudo”-test set $\mathcal{D}_{\text{test}}$. For each $(X_{\operatorname{test}},Y_{\operatorname{test}})\in\mathcal{D}_{\text{test}}$, the transformer output $M_{\theta}(X_{\text{test}},\mathcal{D}_{\text{train}})$ is an approximation to the PPD $p(-|X_{\operatorname{test}},\mathcal{D}_{\text{train}})$. The loss
@@
can be showed to be an unbiased estimate of the conditional KL divergence between the true and approximate PPDs, and is hence minimized via stochastic gradient descent (SGD). We refer the reader to Hollmann et al. [20] for further details on the prior as well as on model training.
@@ 3 Case studies
3.1 Case study I: parameter estimation in a semi-supervised setting
3.1.1 Problem and motivation.
@@ 3.1 Case study I: parameter estimation in a semi-supervised setting
  • A labeled dataset ${\mathcal{D}}_{L}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ containing $n$ IID observations from the joint distribution $P_{X,Y}$;
  • An unlabeled dataset ${\mathcal{D}}_{U}=\{X_{i}\}_{i=n+1}^{n+m}$ containing $m$ IID observations from the marginal distribution $P_{X}$.
@@
$\theta^{*}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{P_{X,Y}}\{l(\theta;X,Y)\},$ (1)
@@
  • Angelopoulos et al. [1] proposed a simple imputation and debiasing framework that makes use of a black-box regression or classification model $M$: (a) Use $M$ to impute labels $\hat{Y}_{i}$ on both the labeled and unlabeled datasets. Denote these imputed datasets as $\widehat{{\mathcal{D}}}_{L}=\{(X_{i},\widehat{Y}_{i})\}_{i=1}^{n}$ and $\widehat{{\mathcal{D}}}_{U}=\{(X_{i},\widehat{Y}_{i})\}_{i=n+1}^{n+m}$ respectively; (b) Estimate the imputation bias via: $\Delta=\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{L})-\widehat{\theta}_{\text{ERM}}({\mathcal{D}}_{L})$; (c) Return the estimate $\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{U})-\Delta$.
@@
  • TabPFN-Debiased (TabPFN-D): Follow Angelopoulos et al. [1]’s approach, we impute each label using the posterior predictive mean, with PPD approximated by TabPFN supplied with ${\mathcal{D}}_{L}$ as a training set. However, return the estimate $\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{L}\cup\widehat{{\mathcal{D}}}_{U})-\Delta$ (i.e. solve ERM on the union of imputed versions of both the labeled and unlabeled datasets). [^2] We found this version to have superior performance to the original estimator of Angelopoulos et al. [1].
@@
We adhere to the original authors’ implementation recommendations for the SLZ method. As a further baseline, we consider the estimator $\widehat{\theta}_{\text{ERM}}({\mathcal{D}}_{L})$ that only makes use of labeled data (Vanilla).
We consider three types of estimands, corresponding to parameters from linear, logistic, and quantile regression working models (Table 1). [^3] These are called “working models” because the true data-generating process differs from the estimation models. We vary:
@@
3.1.4 Results.
Representative results of our experiments are shown in Figure 1, with the complete set of results deferred to Appendix A. The top left, top right, and bottom left panels show MSE with respect to linear, logistic, and quantile regression respectively when $p=5$, with various combinations of $(n,m)$. In the bottom right panel, we show how the results for logistic regression vary with $p$, observing that SLZ is only able to outperform the TabPFN approaches when $p=4$. In general, we observe that TabPFN-I and TabPFN-D significantly outperform SLZ in terms of MSE. We also decompose MSE into squared bias (shaded bar) and variance (solid bar) for the four methods and observe that while TabPFN-I has larger bias than TabPFN-D (as well as the other two methods), it has significantly smaller variance, which leads to smaller MSE.

### v2

@@ Abstract
Hollmann et al. (Nature 637 (2025) 319–326) recently introduced TabPFN, a transformer-based deep learning model for regression and classification on tabular data, which they claim “outperforms all previous methods on datasets with up to 10,000 samples by a wide margin, using substantially less training time.” Furthermore, they have called TabPFN a “foundation model” for tabular data, as it can support “data generation, density estimation, learning reusable embeddings and fine-tuning”. In this paper, we provide a tailored explanation of how TabPFN works for a statistics audience, by emphasizing its interpretation as approximate Bayesian inference. We then explore the significance of TabPFN to the field of statistics: We show that an out-of-the-box application of TabPFN can sometimes outperform specialized state-of-the-art methods for semi-supervised parameter estimation, prediction under covariate shift, and heterogeneous treatment effect estimation. As a partial explanation for the predictive effectiveness of TabPFN, we show that it can simultaneously adapt to both nonparametric structure and parametric structure, for instance, sometimes outperforming LASSO even when assumptions are correctly specified. All experiments can be reproduced using the code provided at https://github.com/qinglong-tian/tabpfn_study.
@@ 1 Introduction
Deep learning has achieved remarkable success in modeling unstructured data types such as text and images. However, for tabular data—arguably the most prevalent and consequential data type across science—ensemble tree methods have remained the gold standard predictive modeling strategy, consistently outperforming even deep learning on benchmark datasets [11, 19, 40, 24]. This long-held wisdom has now been challenged by Hollmann et al. [30], who introduced a deep learning method for tabular data, called TabPFN [^1] The version of the TabPFN model used in our study is the one introduced by Hollmann et al. [30], referred to there as TabPFNv2. While our paper was undergoing revision, the developers released an updated model, TabPFN-2.5 [25]. For consistency, all experiments and discussion in this paper are based on TabPFNv2., which they claimed “outperforms all previous methods on datasets with up to 10,000 samples by a wide margin, using substantially less training time.” As evidence, they showed that TabPFN significantly outperforms random forests, XGBoost, LightGBM, and CatBoost on 5050+ benchmark classification and regression datasets.
@@
Beyond its impressive predictive accuracy, TabPFN is claimed to be robust to outliers and missing data, and automatically estimates predictive distributions, thereby providing uncertainty quantification. Furthermore, it is claimed to support “data generation, density estimation, learning reusable embeddings and fine-tuning”. Because of this wide range of capabilities, Hollmann et al. [30] call TabPFN a “tabular foundation model” [52], appropriating a term that has been used to describe the transformative role played by large language models (LLMs) [7] in natural language processing. TabPFN shares other similarities with LLMs: It makes use of the same type of deep learning architecture (transformers) [53] and the same learning paradigm (in-context learning) [9], which differs fundamentally from traditional statistical and machine learning approaches.
@@
Our goal in this paper is not to further test the claims of Hollmann et al. [30], which have now been further validated on other benchmarks and by other sources (see for instance Erickson et al. 18). Instead, we aim to provide an explanation of how TabPFN works that is tailored to a statistics audience (Section 2) and, more importantly, to explore the significance of TabPFN to the field of statistics. Towards this goal, we first demonstrate the utility of TabPFN in handling three currently researched estimation problems, namely semi-supervised parameter estimation (Section 3.1), heterogeneous treatment effects estimation (Section 3.2), and prediction under covariate shift (Section 3.3). For each problem, we compare a simple out-of-the-box application of TabPFN to a “state-of-the-art” method proposed by a selected paper recently published at a top statistics journal. Using the same simulation set-up described by the paper, augmented with further comparisons on real datasets, we show that TabPFN often achieves better results than the comparator method.
In addition, we investigate how the statistical principle of adaptivity [16] can provide a partial explanation for the predictive effectiveness of TabPFN. In classical nonparametric statistics, adaptivity refers to a procedure’s ability to automatically adjust to unknown aspects of the data-generating process, such as smoothness, sparsity, or latent structure, achieving optimal rates without relying on these properties being specified in advance. In Section 4, we perform numerical experiments suggesting that TabPFN, without any hyperparameter tuning, can simultaneously adapt to both parametric and nonparametric structure: It is competitive with and sometimes even outperforms parametric methods (e.g., LASSO and linear discriminant analysis) even in settings where the parametric assumptions are correct. At the same time, it outperforms local averaging and smoothing methods in misspecified settings—for example, classification problems with substantial label noise. Such adaptivity properties are beyond the reach of ensemble tree methods and other previous machine learning or nonparametric methods.
While TabPFN is not a perfect model, its capabilities indicate that “in-context learning” with a “foundation model” could become a powerful alternative to conventional statistical methodologies across a variety of tasks. This emerging paradigm appears sufficiently promising to merit serious consideration alongside more established approaches. We define these terms and elaborate on their implications in Section 5, where we also discuss avenues for statisticians to guide and influence the field’s development.
@@ 2 An overview of TabPFN
TabPFN makes use of in-context learning (ICL), which is a new learning paradigm that differs from the standard statistical learning approach to regression and classification and has gathered significant interest in the machine learning community. The meaning of ICL has changed over time. In TabPFN, ICL takes the form of approximate Bayesian inference; we elaborate on this perspective below and defer a discussion on the historical development of ICL to Section 5.
The first key idea of TabPFN is to approximate posterior predictive distributions (PPD) using a transformer model, which is the same deep learning architecture that powers LLMs. More precisely, let $\Pi$ denote the space of joint distributions on covariate-label pairs $(X,Y)$ and let $n$ be a fixed sample size. Let $p$ be a prior on $\Pi$ and an observed training data $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is drawn independently identically distributed (IID) from some fixed distribution $\pi^{*}\in\Pi$, let $p(\pi|\mathcal{D})$ denote the posterior on $\Pi$ given $\mathcal{D}$. With slight abuse of notation, the posterior predictive distribution (PPD) for $Y_{n+1}$ given $X_{n+1}=x$ is defined as
$p(y|x,\mathcal{D})=\int_{\Pi}\pi(y|x)p(\pi|\mathcal{D})d\pi.$
@@
Approximating the PPD of a Bayesian regression or classification model is far from a new idea. Indeed, the popular Bayesian Additive Regression Trees (BART) algorithm does the same job via Markov chain Monte Carlo (MCMC) [13]. The novelty of the PFN approach is that while previous approaches (MCMC, variational inference, likelihood-free inference, etc.) attempt to approximate the map $x\mapsto p(y|x,\mathcal{D})$ for a fixed training dataset $\mathcal{D}$, PFN tries to approximate the map jointly over $x$ and $\mathcal{D}$. It has the computational benefit of amortizing computation, which is the basis of the “substantially less training time” claim by Hollmann et al. [30] regarding TabPFN. [^2] Similar ideas have been explored in the variational inference community under the name of Neural Processes (NP) [23]. Both the NP and PFN approaches use a “meta-learning” perspective to approximate the PPD, but they differ in their implementation choices. NPs employ an encoder-decoder architecture with a parametric distribution over the latent variable $z$, naturally lending itself to variational inference. PFNs, on the other hand, adopt a transformer architecture that avoids explicit distributional assumptions in its internal representations. As a result, their training objectives differ slightly: NPs explicitly optimize a variational lower bound, while PFNs frame learning as a direct optimization problem. Such an approximation may also yield statistical benefits via implicit regularization that are not yet fully understood.
@@
The second key idea of TabPFN is its choice of prior on the space of joint distribution $\Pi$. The large number of parameters of a transformer model ($\approx$ 7M in the case of TabPFN) means that it could potentially approximate PPDs defined in terms of complicated priors. TabPFN chooses to define its prior in terms of structural causal models (SCMs) [41]. An SCM defines a directed acyclic graph $\mathcal{G}$ whose nodes represent variables $Z_{1},\ldots,Z_{p}$ and whose edges represent causal relationships. Specifically, each variable $Z_{k}$ is defined via the equation $Z_{k}=f_{k}(Z_{\operatorname{PA}_{\mathcal{G}}(k)},\varepsilon_{k})$, where $f_{k}$ is a deterministic function, $\operatorname{PA}_{\mathcal{G}}(k)$ denotes the parents of $Z_{k}$, and $\varepsilon_{1},\ldots,\varepsilon_{p}$ are independent “exogenous” random variables. SCMs have long been used to model causal inference from observational data, in which case $f_{1},\ldots,f_{p}$ are usually defined to be linear functions. Instead, TabPFN allows them to be neural networks or decision trees, thereby modeling nonlinear and even discontinuous structure.
@@
These two key ideas cohere to provide an efficient training strategy for TabPFN. To fit the model’s parameters $\theta$, a collection $\mathfrak{D}$ of approximately 130M synthetic datasets were generated. Each dataset $\mathcal{D}\in\mathfrak{D}$ contains IID examples from a joint distribution $\pi$, which is itself drawn uniformly from the SCM prior. Note that while $\pi$ may have a complex density, the SCM structure gives a simple way to sample from it.
Next, $\mathcal{D}$ is partitioned into a “pseudo”-training set $\mathcal{D}_{\text{train}}$ and “pseudo”-test set $\mathcal{D}_{\text{test}}$. For each $(X_{\operatorname{test}},Y_{\operatorname{test}})\in\mathcal{D}_{\text{test}}$, the transformer output $M_{\theta}(X_{\text{test}},\mathcal{D}_{\text{train}})$ approximates the PPD $p(\cdot|X_{\operatorname{test}},\mathcal{D}_{\text{train}})$. The loss
@@
can be showed to be an unbiased estimate of the conditional KL divergence between the true and approximate PPDs, and is minimized via stochastic gradient descent. We refer the reader to Hollmann et al. [30] for further details on the prior as well as on model training.
@@ 3 Case studies
3.1 Study I: parameter estimation in a semi-supervised setting
3.1.1 Problem and motivation
@@
  • A labeled dataset ${\mathcal{D}}_{L}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ with $n$ IID observations from the joint $P_{X,Y}$;
  • An unlabeled dataset ${\mathcal{D}}_{U}=\{X_{i}\}_{i=n+1}^{n+m}$ with $m$ IID observations from the marginal $P_{X}$.
@@
$\theta^{*}:=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{P_{X,Y}}\{l(\theta;X,Y)\},$ (1)
@@ 3.1.2 State-of-the-art
  • Angelopoulos et al. [1] proposed a simple imputation and debiasing framework that makes use of an existing black-box regression or classification model $M$: (a) Use $M$ to impute labels $\hat{Y}_{i}$ on both the labeled and unlabeled datasets. Denote these imputed datasets as $\widehat{{\mathcal{D}}}_{L}=\{(X_{i},\widehat{Y}_{i})\}_{i=1}^{n}$ and $\widehat{{\mathcal{D}}}_{U}=\{(X_{i},\widehat{Y}_{i})\}_{i=n+1}^{n+m}$ respectively; (b) Estimate the imputation bias via: $\Delta=\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{L})-\widehat{\theta}_{\text{ERM}}({\mathcal{D}}_{L})$; (c) Return the estimate $\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{U})-\Delta$.
@@ 3.1.3 Experimental design
  • TabPFN-Debiased (TabPFN-D): Following Angelopoulos et al. [1]’s approach, we impute each label using the posterior predictive mean (for continuous labels) or mode (for discrete labels), with PPD approximated by TabPFN supplied with ${\mathcal{D}}_{L}$ as a training set. [^3] Angelopoulos et al. [1] do not specify how to fit the imputation model. However, return the estimate $\widehat{\theta}_{\text{ERM}}(\widehat{{\mathcal{D}}}_{L}\cup\widehat{{\mathcal{D}}}_{U})-\Delta$ (i.e. solve ERM on the union of imputed versions of both the labeled and unlabeled datasets).
@@
We use the original authors’ implementation for the SLZ method. As a further baseline, we consider the estimator $\widehat{\theta}_{\text{ERM}}({\mathcal{D}}_{L})$ that only makes use of labeled data (Vanilla).
We consider three types of estimands, corresponding to parameters from linear, logistic, and quantile regression working models (Table 1). These are called “working models” because the true data-generating process differs from the estimation models. We vary:
@@
3.1.4 Results
Representative results of our experiments are shown in Figure 1, with the complete set of results deferred to Appendix A. The first three panels (from the left) show MSE with respect to linear, logistic, and quantile regression respectively when $p=5$, with various combinations of $(n,m)$. In the last panel, we show how the results for logistic regression vary with $p$, observing that SLZ is only able to outperform the TabPFN approaches when $p=4$. In general, we observe that TabPFN-I and TabPFN-D significantly outperform SLZ in terms of MSE. We also decompose MSE into squared bias (shaded bar) and variance (solid bar) for the four methods and observe that while TabPFN-I has larger bias than TabPFN-D (as well as other two methods), it has significantly smaller variance, which offsets larger bias and leads to smaller MSE.